OpenStudy (safia21):

∆ABC if BC is one inch longer AB, AC is 10 inches shorter than the sum of BC and and AB and the perimeter of ∆ABC is 72 inches find the length of BC

OpenStudy (anonymous):

let AB be x BC is one inch longer AB so,BC = x+1 AC is 10 inches shorter than the sum of BC and and AB so,AC = AB+BC-10 = x+x+1-10 perimeter=72 AB+BC+AC=72 x+x+1+x+x+1-10=72 4x-8=72 4x=80 x=80/4 x=20 BC=x+1 BC=20+1 BC=21